OpenStudy (anonymous):

f(x) = 3x + 2; g(x) = 3x - 5 Find f/g.

OpenStudy (anonymous):

\[(f/g)(x) =\frac{ 3x-5 }{ 3x+2 } ; domain {x|x ≠ - \frac{ 2 }{ 3 }}\] \[ (f/g)(x) =\frac{ 3x+2 }{ 3x-5 } ; domain {x|x ≠\frac{ 5 }{ 3 } }\] \[(f/g)(x) =\frac{ 3x-5 }{ 3x+2 } ; domain {x|x ≠ - \frac{ 5 }{ 3 }}\] \[ (f/g)(x) =\frac{ 3x+2 }{ 3x-5 } ; domain {x|x ≠\frac{ 2 }{ 3 } }\]

OpenStudy (anonymous):

ok what is f?

OpenStudy (anonymous):

3x+2

OpenStudy (anonymous):

i'm guessing the answer is B, or D if i'm right how do i find the domain???

OpenStudy (anonymous):

chose your mind B or D

OpenStudy (anonymous):

D???

OpenStudy (anonymous):

ok plug in x=2/3 what do you get?

OpenStudy (anonymous):

1.66 or 2/3